# Query complexities of quantum channel discrimination and estimation: A unified approach

### v1

@@ 1.2 Summary of results
To summarize our results succinctly, we first either recall or establish upper bounds on the squared Bures distance in both the parallel and adaptive settings of channel discrimination (see Theorems 16 and 18), and we do the same for the symmetric logarithmic derivative (SLD) Fisher information for channel estimation (see Theorems 21 and 22). After doing so, we then establish lower bounds on the error probabilities of channel discrimination and estimation in terms of the squared Bures distance and SLD Fisher information (see Corollaries 24, 30, and 33), following a line of thinking adopted in [CDL+25]. By applying the definitions of query complexities for channel estimation and discrimination, we then convert these lower bounds on error probabilities to lower bounds on the query complexities for these tasks (see Corollaries 25, 31, and 33). Finally, we provide various numerical methods for efficiently computing lower bounds on the query complexities of channel discrimination and estimation, in both the parallel and adaptive access models.
@@
Note on independent work—While finalizing our manuscript, we noticed that the concurrent and independent paper [SDD25] appeared. Similar to the theme of our paper and that of [YF17], Ref. [SDD25] presents a unified approach to quantum channel discrimination and estimation. More specifically, the bound presented in Eqs. (18)–(20) therein are similar in spirit to our bound from Corollary 24.
@@ 1.3 Paper organization
  • Section 5 reviews the SLD Fisher information of states and channels. In particular, Theorems 11 and 12 by and large date back to [FI08, Theorems 1 and 4]. However, some of the alternative expressions for SLD Fisher information of states, as presented in Theorem 11, appear to be novel, and our proof of Theorem 11 is distinct from the proof of [FI08, Theorem 1]. A similar statement applies to Theorem 12.
@@
  • Section 7 reviews upper bounds on the SLD Fisher information of channels in the parallel and adaptive settings of channel estimation. The upper bound in the parallel setting (Theorem 21) was already established in [FI08, Section 4.1], and the upper bound in the adaptive setting (Theorem 22) was established somewhat recently in [KGADD23, Eq. (11)], improving upon an earlier bound from [DDM14, Eq. (9)]. Interestingly, our proofs for Theorems 21 and 22 mirror those given for Theorems 16 and 18, respectively, making use of basic properties of isometries and spectral norms, again consistent with our overall theme. Furthermore, our proofs for Theorems 16 and 18 can be understood as replacing derivatives with finite differences, when compared to the proofs for Theorems 21 and 22, consistent with our other theme of having a unified approach to channel discrimination and estimation.
  • Section 8 presents our main results: lower bounds on the error probabilities and query complexities of channel discrimination and estimation. Ultimately, with the framework in place at this point, the results of this section follow somewhat directly from definitions, reasoning similar to that in [CDL+25, NW25], and the bounds from Theorems 16, 18, 21, and 22.
@@ 3.1.1 Parallel and adaptive settings of channel discrimination
The first inequality follows because a parallel channel discrimination strategy is a special case of an adaptive channel discrimination strategy. The second inequality is a consequence of [ACMT+07, Theorem 1], as remarked in [CDL+25, Eqs. (D1)–(D3)], and the fact that $\min\left\{p,q\right\}=\min_{s\in\left[0,1\right]}p^{s}q^{1-s}$. As such, the following inequalities hold for $p\in\left(0,1\right)$:
@@ 3.1.2 Query complexities of channel discrimination
$n_{\|}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\|}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\leq\varepsilon\right\},$ (3.9)
$n_{\operatorname{ad}}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\operatorname{ad}}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\leq\varepsilon\right\}.$ (3.10)
@@
$n_{\operatorname{ad}}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)=n_{\|}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)=1.$ (3.12)
@@ 3.2.1 Parallel setting
Let us begin by considering the parallel setting of channel estimation. An $n$-query parallel estimation protocol begins with the estimator preparing a state $\rho_{RA_{1}\cdots A_{n}}$, feeding the systems $A_{1}\cdots A_{n}$ of this state through the unknown channel $\mathcal{N}_{\theta}^{\otimes n}$, and performing a measurement $\smash{\mathcal{Q}^{\left(n\right)}\coloneqq(Q_{\hat{\theta}}^{\left(n\right)})_{\hat{\theta}\in\Theta}}$ on the reference system $R$ and the channel output systems $B_{1}\cdots B_{n}$. The probability of outputting the estimate $\hat{\theta}$, given that $\theta$ is the true parameter value, is as follows:
@@
$n_{\|}^{*}(\varepsilon,\delta,\mathcal{F})\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\|}(\delta,\mathcal{F},n)\leq\varepsilon\right\}.$ (3.21)
@@ 3.2.2 Adaptive setting
$n_{\operatorname{ad}}^{*}(\varepsilon,\delta,\mathcal{F})\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)\leq\varepsilon\right\}.$ (3.29)
@@ 3.3 Connecting channel estimation and discrimination
$p_{e,\|}(\delta,\mathcal{F},n)$ $\geq\sup_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\\ \theta,\theta^{\prime}\in\Theta\end{subarray}}p_{e,\|}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},n\right),$ (3.30)
@@
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)$ $\geq\sup_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\\ \theta,\theta^{\prime}\in\Theta\end{subarray}}p_{e,\operatorname{ad}}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},n\right).$ (3.31)
@@
That is, the minimax error probabilities for channel estimation, with tolerance $\delta$, are not smaller than the channel discrimination error probabilities for two channels with parameter values that are more than $2\delta$-far from each other. Stated informally, channel estimation is just as difficult as channel discrimination of nearby channels. As such, limitations on channel discrimination provide limitations on channel estimation.
@@
$n_{\|}^{*}(\varepsilon,\delta,\mathcal{F})$ $\geq\sup_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\\ \theta,\theta^{\prime}\in\Theta\end{subarray}}n_{\|}^{*}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},\varepsilon\right),$ (3.32)
$n_{\operatorname{ad}}^{*}(\varepsilon,\delta,\mathcal{F})$ $\geq\sup_{\begin{subarray}{c}\left|\theta-\theta^{\prime}\right|>2\delta,\\ \theta,\theta^{\prime}\in\Theta\end{subarray}}n_{\operatorname{ad}}^{*}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},\varepsilon\right).$ (3.33)
@@ 4.1.1 Fidelity of states
$\sqrt{F}(\rho,\sigma)$ $=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right]$ (4.4)
@@
$\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}$ $=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\left|\operatorname{Tr}\!\left[\sqrt{\rho}\sqrt{\sigma}W\right]\right|$ (4.8)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\operatorname{Tr}\!\left[\sqrt{\rho}\sqrt{\sigma}W\right]\right]$ (4.9)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\langle\Gamma|\left(I\otimes\sqrt{\rho}\sqrt{\sigma}W\right)|\Gamma\rangle\right]$ (4.10)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\langle\Gamma|\left(W^{T}\otimes\sqrt{\rho}\sqrt{\sigma}\right)|\Gamma\rangle\right]$ (4.11)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\langle\Gamma|\left(I\otimes\sqrt{\rho}\right)\left(W^{T}\otimes I\right)\left(I\otimes\sqrt{\sigma}\right)|\Gamma\rangle\right]$ (4.12)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\langle\psi^{\rho}|\left(W^{T}\otimes I\right)|\psi^{\sigma}\rangle\right]$ (4.13)
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right].$ (4.14)
@@
$=\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right].$ (4.22)

### v2

@@ 1.2 Summary of results
To summarize our results succinctly, we first either recall or establish upper bounds on the squared Bures distance in both the parallel and adaptive settings of channel discrimination (see Theorems 17 and 19), and we do the same for the symmetric logarithmic derivative (SLD) Fisher information for channel estimation (see Theorems 22 and 23). After doing so, we then establish lower bounds on the error probabilities of channel discrimination and estimation in terms of the squared Bures distance and SLD Fisher information (see Corollaries 25, 31, and 34), following a line of thinking adopted in [CDL+25]. By applying the definitions of query complexities for channel estimation and discrimination, we then convert these lower bounds on error probabilities to lower bounds on the query complexities for these tasks (see Corollaries 26, 32, and 34). Finally, we provide various numerical methods for efficiently computing lower bounds on the query complexities of channel discrimination and estimation, in both the parallel and adaptive access models.
@@
Note on independent work—While finalizing our manuscript, we noticed that the concurrent and independent paper [SDD26] appeared. Similar to the theme of our paper and that of [YF17], Ref. [SDD26] presents a unified approach to quantum channel discrimination and estimation. More specifically, the bound presented in Eqs. (18)–(20) therein are similar in spirit to our bound from Corollary 25.
@@ 1.3 Paper organization
  • Section 5 reviews the SLD Fisher information of states and channels. In particular, Theorems 12 and 13 by and large date back to [FI08, Theorems 1 and 4]. However, some of the alternative expressions for SLD Fisher information of states, as presented in Theorem 12, appear to be novel, and our proof of Theorem 12 is distinct from the proof of [FI08, Theorem 1]. A similar statement applies to Theorem 13.
@@
  • Section 7 reviews upper bounds on the SLD Fisher information of channels in the parallel and adaptive settings of channel estimation. The upper bound in the parallel setting (Theorem 22) was already established in [FI08, Section 4.1], and the upper bound in the adaptive setting (Theorem 23) was established somewhat recently in [KGADD23, Eq. (11)], improving upon an earlier bound from [DDM14, Eq. (9)]. Interestingly, our proofs for Theorems 22 and 23 mirror those given for Theorems 17 and 19, respectively, making use of basic properties of isometries and spectral norms, again consistent with our overall theme. Furthermore, our proofs for Theorems 17 and 19 can be understood as replacing derivatives with finite differences, when compared to the proofs for Theorems 22 and 23, consistent with our other theme of having a unified approach to channel discrimination and estimation.
  • Section 8 presents our main results: lower bounds on the error probabilities and query complexities of channel discrimination and estimation. Ultimately, with the framework in place at this point, the results of this section follow somewhat directly from definitions, reasoning similar to that in [CDL+25, NW25], and the bounds from Theorems 17, 19, 22, and 23.
@@ 2 Preliminaries
$\mathbb{B}$ Set of contractions $\mathbb{B}\coloneqq\left\{W\in\mathbb{L}:\left\|W\right\|\leq 1\right\}$
@@ 3.1.1 Parallel and adaptive settings of channel discrimination
The first inequality follows because a parallel channel discrimination strategy is a special case of an adaptive channel discrimination strategy. The second inequality is a consequence of [ACMT+07, Theorem 1], as remarked in [CDL+25, Eqs. (D1)–(D3)], and the fact that $\min\!\left\{p,q\right\}=\min_{s\in\left[0,1\right]}p^{s}q^{1-s}$. Alternatively, one can understand this inequality as resulting from discarding the channel outputs and simply guessing the channel that has the higher prior probability $p$ or $q$. Discarding the channel in either (3.1) or (3.4) leads to the following expression for the minimum error probability $\frac{1}{2}\left(1-\left|p-q\right|\right)$, which is equal to $\min\{p,q\}$. As such, the following inequalities hold for $p\in\left(0,1\right)$:
@@ 3.1.2 Query complexities of channel discrimination
$n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\|}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\leq\varepsilon\right\},$ (3.9)
$n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\operatorname{ad}}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},n)\leq\varepsilon\right\}.$ (3.10)
@@
$n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)=n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)=1.$ (3.12)
@@ 3.2.1 Parallel setting
Let us begin by considering the parallel setting of channel estimation. An $n$-query parallel estimation protocol begins with the estimator preparing a state $\rho_{RA_{1}\cdots A_{n}}$, feeding the systems $A_{1}\cdots A_{n}$ of this state through the unknown channel $\mathcal{N}_{\theta}^{\otimes n}$, and performing a measurement $\smash{\mathcal{Q}^{\left(n\right)}\coloneqq(Q_{\hat{\theta}}^{\left(n\right)})_{\hat{\theta}\in\Theta}}$ on the reference system $R$ and the channel output systems $B_{1}\cdots B_{n}$. We assume that $\mathcal{Q}^{\left(n\right)}$ forms a positive operator-valued measure (POVM) on $\Theta$ with respect to the Lebesgue measure. The probability of outputting the estimate $\hat{\theta}$, given that $\theta$ is the true parameter value, is as follows:
@@
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\|}(\delta,\mathcal{F},n)\leq\varepsilon\right\}.$ (3.21)
@@ 3.2.2 Adaptive setting
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})\coloneqq\inf\left\{n\in\mathbb{N}:p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)\leq\varepsilon\right\}.$ (3.29)
@@ 3.3 Connecting channel estimation and discrimination
Lemma 7.
For all $\delta>0$ and $n\in\mathbb{N}$ and for every channel family $\mathcal{F}$, the following bounds hold:
@@
$p_{e,\|}(\delta,\mathcal{F},n)$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}p_{e,\|}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},n\right),$ (3.30)
@@
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}p_{e,\operatorname{ad}}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},n\right).$ (3.31)
Proof.
We provide a simple proof of (3.31) for completeness, which follows Le Cam’s two-point method [Cam73, Yu97]. The proof of (3.30) is analogous. Define the measurement operator $Q_{\left[\theta-\delta,\theta+\delta\right]}^{\left(n\right)}$ as follows:
$Q_{\left[\theta-\delta,\theta+\delta\right]}^{\left(n\right)}\coloneqq\int_{\left|\hat{\theta}-\theta\right|\leq\delta}d\hat{\theta}\>Q_{\hat{\theta}}^{\left(n\right)}.$ (3.32)
Fix $\theta_{1},\theta_{2}\in\Theta$ such that $\left|\theta_{1}-\theta_{2}\right|>2\delta$. Then
$p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)$
$=\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\theta\in\Theta}\left\{\int_{\hat{\theta}:\left|\hat{\theta}-\theta\right|>\delta}d\hat{\theta}\>\operatorname{Tr}\!\left[Q_{\hat{\theta}}^{\left(n\right)}\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.33)
$=\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\theta\in\Theta}\left\{\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta-\delta,\theta+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.34)
$\geq\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\begin{subarray}{c}\theta_{1},\theta_{2}\in\Theta:\\ |\theta_{1}-\theta_{2}|>2\delta\end{subarray}}\left\{\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{1}}^{(n)}\!\left(\rho_{RA}\right)\right],\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{2}-\delta,\theta_{2}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{2}}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.35)
$\geq\inf_{\mathcal{E}^{\left(n\right)}}\sup_{\begin{subarray}{c}\theta_{1},\theta_{2}\in\Theta:\\ |\theta_{1}-\theta_{2}|>2\delta\end{subarray}}\left\{\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{1}}^{(n)}\!\left(\rho_{RA}\right)\right],\operatorname{Tr}\!\left[Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\mathcal{A}_{\theta_{2}}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.36)
$\geq\inf_{\mathcal{E}^{\left(n\right)}}\left\{\frac{1}{2}\operatorname{Tr}\!\left[\left(I-Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\right)\mathcal{A}_{\theta_{1}}^{(n)}\!\left(\rho_{RA}\right)\right]+\frac{1}{2}\operatorname{Tr}\!\left[Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}\mathcal{A}_{\theta_{2}}^{(n)}\!\left(\rho_{RA}\right)\right]\right\}$ (3.37)
$=p_{e,\operatorname{ad}}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{1}},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{2}},n\right).$ (3.38)
The second inequality follows because $I-Q_{\left[\theta_{2}-\delta,\theta_{2}+\delta\right]}^{\left(n\right)}\geq Q_{\left[\theta_{1}-\delta,\theta_{1}+\delta\right]}^{\left(n\right)}$, given that $\left|\theta_{1}-\theta_{2}\right|>2\delta$. Indeed, since $|\theta_{1}-\theta_{2}|>2\delta$, the intervals $\theta_{1}-\delta,\theta_{1}+\delta]$ and $\theta_{2}-\delta,\theta_{2}+\delta]$ are disjoint. Because $\big(Q_{\hat{\theta}}^{(n)}\big)_{\hat{\theta}\in\Theta}$ forms a POVM, it is countably additive on disjoint measurable sets. Hence,
$Q_{[\theta_{1}-\delta,\theta_{1}+\delta]}^{(n)}+Q_{[\theta_{2}-\delta,\theta_{2}+\delta]}^{(n)}=Q_{[\theta_{1}-\delta,\theta_{1}+\delta]\cup[\theta_{2}-\delta,\theta_{2}+\delta]}^{(n)}\leq I,$ (3.39)
where $I$ denotes the identity operator on the final output space. Therefore,
$Q_{[\theta_{1}-\delta,\theta_{1}+\delta]}^{(n)}\leq I-Q_{[\theta_{2}-\delta,\theta_{2}+\delta]}^{(n)},$ (3.40)
which proves the claimed operator inequality. The last equality follows because the optimization ranges over all adaptive discrimination strategies between $\mathcal{N}_{\theta_{1}}$ and $\mathcal{N}_{\theta_{2}}$ with equal prior probabilities, given that every adaptive estimation protocol allows for an arbitrary final POVM on the output system. Hence, the expression coincides with the minimum Bayesian error probability for adaptive discrimination between these two channels. Since the inequality $p_{e,\operatorname{ad}}(\delta,\mathcal{F},n)\geq p_{e,\operatorname{ad}}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{1}},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta_{2}},n\right)$ holds for all $\theta_{1},\theta_{2}\in\Theta$ such that $\left|\theta_{1}-\theta_{2}\right|>2\delta$, we conclude (3.31). ∎
@@
Lemma 7 states that the minimax error probabilities for channel estimation, with tolerance $\delta$, are not smaller than the channel discrimination error probabilities for two channels with parameter values that are more than $2\delta$-far from each other. Stated informally, channel estimation is just as difficult as channel discrimination of nearby channels. As such, limitations on channel discrimination provide limitations on channel estimation.
@@
$n_{\|}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}n_{\|}^{\star}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},\varepsilon\right),$ (3.41)
$n_{\operatorname{ad}}^{\star}(\varepsilon,\delta,\mathcal{F})$ $\geq\sup_{\begin{subarray}{c}\theta,\theta^{\prime}\in\Theta:\\ \left|\theta-\theta^{\prime}\right|>2\delta\end{subarray}}n_{\operatorname{ad}}^{\star}\!\left(\nicefrac{{1}}{{2}},\mathcal{N}_{\theta},\nicefrac{{1}}{{2}},\mathcal{N}_{\theta^{\prime}},\varepsilon\right).$ (3.42)
@@ 4.1.1 Fidelity of states
$\sqrt{F}(\rho,\sigma)$ $=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right]$ (4.4)
@@
$\left\|\sqrt{\rho}\sqrt{\sigma}\right\|_{1}$ $=\sup_{W\in\mathbb{B}}\left|\operatorname{Tr}\!\left[\sqrt{\rho}\sqrt{\sigma}W\right]\right|$ (4.8)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\operatorname{Tr}\!\left[\sqrt{\rho}\sqrt{\sigma}W\right]\right]$ (4.9)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\Gamma|\left(I\otimes\sqrt{\rho}\sqrt{\sigma}W\right)|\Gamma\rangle\right]$ (4.10)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\Gamma|\left(W^{T}\otimes\sqrt{\rho}\sqrt{\sigma}\right)|\Gamma\rangle\right]$ (4.11)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\Gamma|\left(I\otimes\sqrt{\rho}\right)\left(W^{T}\otimes I\right)\left(I\otimes\sqrt{\sigma}\right)|\Gamma\rangle\right]$ (4.12)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W^{T}\otimes I\right)|\psi^{\sigma}\rangle\right]$ (4.13)
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right].$ (4.14)
@@
$=\sup_{W\in\mathbb{B}}\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right].$ (4.22)